OpenStudy (anonymous):

Need Help!! Solve the system of equations. 8y+10=6x 8y-4x=-12

OpenStudy (solomonzelman):

subtract the second equation from the first.

OpenStudy (solomonzelman):

But before you do this, you add 4x to both sides add 12 to both sides

OpenStudy (anonymous):

so it would be like 8y+10=6x -8y+12=+4x

OpenStudy (solomonzelman):

NO

OpenStudy (solomonzelman):

it is positive 8y in the 2nd equation.

OpenStudy (solomonzelman):

you just added 12, and then 4x to both sides, 8y doesn't change.

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

typed something extra lol

OpenStudy (solomonzelman):

Ok, so what are your equations are like?

OpenStudy (anonymous):

will it be 14y+22=10x

OpenStudy (solomonzelman):

No, hoe do you get that, you have so far, 8y+10=6x 8y+12=4x

OpenStudy (imtiaz7):

8y+10=6x > 8y-6x=-10 8y-4x=-12 > 8y-4x=-12 Solution: Subtract above both equations: we get: -2x=+2 x=-1 put x's value in eq. 2 that is 8y-4x=-12 so we get: 8y-4x=-12 8y-4(-1)=-12 8y+4=-12 8y=-12-4 8y=-16 y=-16/8 y=-2 so we get the values x=-1 and y=-2

OpenStudy (anonymous):

so i add them no?

OpenStudy (solomonzelman):

Subtract them from each other. 8y+10=6x - 8y+12=4x ---------

OpenStudy (anonymous):

thanks @Imtiaz7 but i want to know how to do it step by step

OpenStudy (anonymous):

-2=2x

OpenStudy (solomonzelman):

So subtract them from each other. 8y+10=6x - 8y+12=4x --------- 0x - 2 = 2x -2=2x x=?

OpenStudy (anonymous):

x=-1

OpenStudy (solomonzelman):

yup!

OpenStudy (anonymous):

so i put this with the other equation

OpenStudy (solomonzelman):

Plug -1 in instead of x, into any of the equations to solve for y.

OpenStudy (solomonzelman):

Yes, GREAT!

OpenStudy (anonymous):

y=-2

OpenStudy (imtiaz7):

ok i want to tell the steps that we first look out these two equation they were solve by adding or substraction or by substitution so here is 8y in both equations and also they are positive so we subtracted both equation to remove the one veriable to get the value of other so 8y are cancelled and soling these we get the value of x, by substituting the x's value in any one of these two eq. we got the value of y. this is all that i did for u

OpenStudy (anonymous):

thank you guys

OpenStudy (imtiaz7):

ok(thank you guys)...

OpenStudy (solomonzelman):

YUP!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

If you want to ask a new question, please open a new thread.

OpenStudy (anonymous):

no its about the same question

OpenStudy (solomonzelman):

Oh, OK, go.

OpenStudy (solomonzelman):

So you had x=-1 y=-2 And I see you followed the steps.

OpenStudy (anonymous):

what would be the resulting equation if we eliminate the y variable?

OpenStudy (solomonzelman):

8y+10=6x - 8y+12=4x ---------- 0y -2=2y -2=2y

OpenStudy (anonymous):

but wouldnt it end up being y=(answer)

OpenStudy (solomonzelman):

well, y is a part of the answer, you are looking for x and y.

OpenStudy (solomonzelman):

it would certainly lead to finding y, if you divide both sides by -2.

OpenStudy (imtiaz7):